Consider the differential equation e24 y – 4y +4y= t> 0. t2 (a) Find T1, T2, roots of the characteristic polynomial of the eConsider the differential equation e24 y – 4y +4y= t> 0. t2 (a) Find T1, T2, roots of the characteristic polynomial of the eConsider the differential equation 5e2t y – 4y + 4y = t>0. t2 (a) Find T1, T2, roots of the characteristic polynomial of the
Answer #1

o We have (D² - 4D+4) y = e²+ 1 +² - 0 a) Put D= m in honwgenious port of - m-4m+4=0 gives (M-2)=0 m=2, 2. Thus soots are 2,-b) Thus, the cof will be (F= (6, + C₂ *) e ²t = C, e24 + (te²t : Fundamental solutions will be -) Y,!!)= e²t 92/8)=4e2t c) W1